# Designing Dynamic Pricing for Bike-sharing Systems via Differentiable Agent-based Simulation

### v1

@@ 3.2 Representation of Trip Demand
[…]
@@ 4.1.1 Scenario Settings
To verify the fundamental characteristics of our method, we generate a simple synthetic trip demand scenario involving two stations: one designed as the high-demand station and the other as the low-demand station (Fig. 4a). The demand are time-invariant, with $\lambda^{dep}_{t,01}=1.0$ and $\lambda^{dep}_{t,10}=0.5$ (Fig. 4b), and $\lambda^{dur}_{t,ij}=1$. Similarly, the discount parameter $p_{t,j}$ is set to be time-invariant. Thus, only two parameters must be estimated: $p_{0}$ and $p_{1}$. Concerning parameters related to the DCM, the weights associated with the differences in discount and distance are set to $w_{\text{discount}}=1$ and $w_{\text{distance}}=-1$, respectively. Here, $asc_{s}$ is set to 0 if station $s$ is the originally intended destination station, as determined by a bicycle departure rate distribution. Conversely, $asc_{s}$ is set to 1-1- 1 if station $s$ is not the originally intended one. The total number of simulation time steps $T$ is set to 100, with each station starting with an initial inventory of 100 bicycles.
@@
[…]
@@
[…]
@@ 4.1.3 Implementation Details
where $Loss$ represents the loss function, and ${I}_{T,j}$ and $\hat{I}_{T,j}$ represent the desired and simulated number of bicycle inventories at station $j$ at the final time step $T$, respectively. All experiments are performed on a CPU with 2$\times$×Xeon Gold 5218 CPU (2.3 GHz $\times$14 core) and 192 GB of RAM.
@@ 4.2.2 Scenario 2: Small-scale Setting
Table 1 lists the $cost$ and loss for each initial discount parameter pattern obtained by each algorithm. As hypothesized in Scenario 1, the results demonstrate that the discount parameters estimated using initial value pattern 1 yield an 76% lower $cost$ than pattern 2, while maintaining comparable loss values. This reveals that setting the initial values for estimating the discount parameters to near zero can reduce the cost associated with user-based relocation using the estimated discount parameters. Although FD also achieves a relatively small $cost$, the loss values do not reach the same level as our method, indicating that it does not obtain a price that can balance the number of bicycle inventories at the final time step.

### v2

@@ 3.2 Representation of Trip Demand
Trip demands is represented using probabilistic distributions that closely approximate actual trip demands in BSS modeling [10, 35]. Specifically, trip demand is modeled using two types of probabilistic distributions: the Poisson distribution for representing the bicycle departure rate (i.e., the timing when the user starts using the bicycle) and the exponential distribution for representing trip duration. These distributions determine the number of potential agents traveling from an origin station $i$ to a destination station $j$ at timestep $t$, with a trip duration of $dur_{t,ij}$. These are characterized by a parameter set for each origin–destination station pair $(i,j)$ at the given timestep $t$. Thus, our ABM has two sets of $J\times J\times T$ parameters for each distribution representing trip demand, as illustrated in Fig 2a. These parameters are described as $\lambda^{dep}_{t,ij}$ and $\lambda^{dur}_{t,ij}$ for the distribution of the bicycle departure rate and trip duration, respectively. According to the definition of the Poisson distribution, $\lambda^{dep}_{t,ij}$ represents the expected value of the number of bicycle departures, corresponding to the number of trips departing from station $i$ to station $j$ at timestep $t$. Therefore, the expected value of the number of trips departing from a certain station $k$ at timestep $t$ is given by $\sum_{j}\lambda^{dep}_{t,kj}$, as illustrated in Fig. 2b.
@@ 4.1.1 Scenario Settings
To verify the fundamental characteristics of our method, we generate a simple synthetic trip demand scenario involving two stations: one designed as the high-demand station and the other as the low-demand station (Fig. 4a). The demand are time-invariant, with $\lambda^{dep}_{t,01}=1.0$ and $\lambda^{dep}_{t,10}=0.5$ (Fig. 4b), and $\lambda^{dur}_{t,ij}=1$. Similarly, the discount parameter $p_{t,j}$ is set to be time-invariant. Thus, only two parameters must be estimated: $p_{0}$ and $p_{1}$. Concerning parameters related to the DCM, the weights associated with the differences in discount and distance are set to $w_{\text{discount}}=1$ and $w_{\text{distance}}=-1$, respectively. Here, $asc_{s}$ is set to 0 if station $s$ is the originally intended destination station, as determined by a bicycle departure rate distribution. Conversely, $asc_{s}$ is set to 1-1 if station $s$ is not the originally intended one. The total number of simulation time steps $T$ is set to 100, with each station starting with an initial inventory of 100 bicycles.
@@
In real-world scenarios within limited spaces such as university campuses, the deployment of BSS involves dozens of stations [10]. As a scenario for deploying BSS at this scale, we assume a simulation period of 10 hours with 0.5 hour time steps (i.e., a total simulation time of $T=20$) across 25 stations, which include both high- and low-demand stations (Fig. 4c), with an initial bicycle inventory of 100. In this scenario, $p_{t,j}$ remains constant for each set of five time steps. Thus, the total number of discount parameters to be estimated is 100 $(25\text{ stations}\times(20\text{ time steps}/5\text{ steps}))$. It is considered that trips over longer distances take more time. To reflect this, the trip duration distribution parameter is designed as $\lambda^{dur}_{t,ij}=5/d_{ij}$. $\lambda^{dep}_{t,ij}$ is designed as a time-variant parameter: for high-demand stations, $\lambda^{dep}_{t,ij}=\exp(-0.2t)+0.1+U(0,0.2)$, and for low-demand stations, $\lambda^{dep}_{t,ij}=0.1+U(0,0.2)$ (Fig.  4d). $U(a,b)$ denotes a uniform distribution between $a$ and $b$. The parameters related to DCM are set to be consistent with those in Scenario 1.
@@
In a large-scale urban deployment of BSS, the number of stations can exceed 100 [7]. To verify our method in large-scale deployments of BSS, we conduct simulations across 289 stations, including both high- and low-demand stations (Fig. 4e). The simulation period, time step, $p_{t,j}$ adjustment, and initial bicycle inventory are consistent with those in Scenario 2. Consequently, the total estimated number of discount parameters is 1156 $(289\text{ stations}\times(20\text{ time steps}/5\text{ steps}))$. As illustrated in Fig. 2b, the expected value of the number of trips departing from each station is proportional to both the number of stations and the parameter $\lambda^{dep}_{t,ij}$. Given that the number of stations is approximately 10 times larger than in Scenario 2, the parameter $\lambda^{dep}_{t,ij}$ is set to be 10 times smaller than in Scenario 2. Specifically, for high-demand stations, $\lambda^{dep}_{t,ij}$ is calculated as $\frac{1}{10}(\exp(-0.2t)+0.1+U(0,0.2))$, and for low-demand stations, it is $\frac{1}{10}(0.1+U(0,0.2))$ (Fig. 4f). This ensures that the number of trips beginning from each station per time step closely matches Scenario 2, enabling experiments under identical inventory conditions. The parameters related to DCM are set to be consistent with those in Scenarios 1 and 2.
@@ 4.1.3 Implementation Details
[…]
@@ 4.2.2 Scenario 2: Small-scale Setting
Table 1 lists the costcost and loss for each initial discount parameter pattern obtained by each algorithm. As hypothesized in Scenario 1, the results demonstrate that the discount parameters estimated using initial value pattern 1 yield an 76% lower costcost than pattern 2, while maintaining comparable loss values. This reveals that setting the initial values for estimating the discount parameters to near zero can reduce the cost associated with user-based relocation using the estimated discount parameters. Although FD also achieves a relatively small costcost, the loss values do not reach the same level as our method, indicating that it does not obtain a price that can balance the number of bicycle inventories at the final time step.